OpenStudy (anonymous):

X A B C D P(X) 0.30 0.20 ? 0.28 find P(A and D) assuming that A and D are independent A. 0.22 B. 0.084 C. 0.02 D. 0.58

OpenStudy (anonymous):

P( A and D) = P(A)*P(D) if A and D are independent.

OpenStudy (anonymous):

This is what I got: P(A n D)=P(A)*P(D) =0.3 times 0.28

OpenStudy (anonymous):

thx